psirockin2:

Can someone help me please? Write an equation in point slope form of the line that goes through the point (9,3) and has a slope of 4/9.

Bob:

y – y1 = m(x – x1) y - 3 = 4/9(x - 9) There's your equation.

Bob:

Just plug them in the point slope formula.

psirockin2:

Okay thanks :)

Bob:

No prob